OpenStudy (anonymous):

Simplify (2x^2-x-15)/(2x+5)

OpenStudy (anonymous):

Use the distributive property or FOIL.

OpenStudy (anonymous):

as I understood: 2x^2-6x-5x-15=2x(x-3)-5(x-3)=(x-3)(2x-5) And there I am lost because of qualms, whwther should I reverse sign(then it will be 2x+5 or to leave it and the answer will be...terrible I'd say)

OpenStudy (anonymous):

Did I help you any Emily9?

OpenStudy (anonymous):

I am still a little bit confused

OpenStudy (anonymous):

Type where you are confused at.

OpenStudy (anonymous):

the converting numerator:2x(x-3)-5(x-3) or 2x(x-3)+5(x-3)?

OpenStudy (anonymous):

I am sorry. This is confusing for me too.